OpenStudy (girlgoyleh.):

Factor this completely? 81x^2-25

ganeshie8 (ganeshie8):

81 and 25 what do they remind you of ?

OpenStudy (girlgoyleh.):

numbers? lol

ganeshie8 (ganeshie8):

Haha They are special numbers, think again

OpenStudy (girlgoyleh.):

they both are composite?

ganeshie8 (ganeshie8):

Good. They are more special than that. Try again

OpenStudy (girlgoyleh.):

they are squared numbers? other than that i dont know :P

ganeshie8 (ganeshie8):

BINGO!

ganeshie8 (ganeshie8):

81 = 9^2 and 25 = 5^2

ganeshie8 (ganeshie8):

Can we write the given expression as \[(9x)^2 - (5)^2\] ?

ganeshie8 (ganeshie8):

Next recall the identity \[a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (girlgoyleh.):

(9x+5)(9x-5)

OpenStudy (skullpatrol):

You got it :-)

OpenStudy (girlgoyleh.):

thanks to the both of you! <3

OpenStudy (skullpatrol):

Thanks for trying to learn :D

ganeshie8 (ganeshie8):

Good job!